Entrance tickets for a carnival are $8 for children and $13 for adults. If one evening 867 tickets were sold with total receipts of $8381, how many children and how many adults entered the carnival?
Answer by rfer(16322)

c+a=867
c=867-a
8c+13a=8381
-------------------
8(867-a)+13a=8381
6936-8a+13a=8381
5a=8381-6936
5a=1445
a=1445/5
a=289 adults
867-289=578 children
------------------
8*578+13*289=
4624+3757=$8381